OpenStudy (anonymous):

Given: f(x) = x2 - 1 and g(x) = 2x + 3. Find (g of f)(2)

OpenStudy (anonymous):

@jim_thompson5910 @Mertsj @Luis_Rivera @hartnn

hartnn (hartnn):

ok, you can first find f(2) =... ? put x=2 in f(x) f(2) = 2^2 -1 =... ?

OpenStudy (anonymous):

That would be 3/2 ?

hartnn (hartnn):

why /2 ? 2^2 = 4 f(2) = 4-1 = 3 right ?

OpenStudy (anonymous):

Ohh, ok

hartnn (hartnn):

so, for (g of f)(2) = g (f(2)) = g (3) now you have to find g(3) just plug in x= 3 in g(x)

OpenStudy (anonymous):

So my answer would be 6?

hartnn (hartnn):

6? g(3) = 2*3 + 3 = ... ?

OpenStudy (anonymous):

Forgot to do plus 3 haha, it's 9

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

yes, 9 welcome ^_^